The Titanic hit an iceberg estimated to be half her mass. Before hitting the iceberg, the Titanic

was estimated to be going 11.3 m/s. What was the momentum of the Titanic?


1
Expert's answer
2022-03-07T18:06:26-0500

The momentum is given as follows:


"p = mv"

where "m\approx 46\times 10^6kg" is the mass of the Titanic, and "v = 11m\/s" is its speed. Thus, obtain:


"p\approx 46\times 10^6kg\cdot 11m/s = 5.06\times 10^8kg\cdot m/s"

Answer. "5.06\times 10^8kg\cdot m/s".